### v2

# Randomized subspace correction methods

Abstract
This paper introduces an abstract framework for randomized subspace correction methods for convex optimization, which unifies and generalizes a broad class of existing algorithms, including domain decomposition, multigrid, and block coordinate descent methods. We provide a convergence rate analysis ranging from minimal assumptions to more practical settings, such as sharpness and strong convexity. While most existing studies on block coordinate descent methods focus on nonoverlapping decompositions and smooth or strongly convex problems, our framework extends to more general settings involving arbitrary space decompositions, inexact local solvers, and problems with weaker smoothness or convexity assumptions. The proposed framework is broadly applicable to convex optimization problems arising in areas such as nonlinear partial differential equations, imaging, and data science.
keywords:
Subspace correction methods, Block coordinate descent methods, Domain decomposition methods, Randomized methods, Convex optimization
{AMS}
90C25, 65N55, 65J05, 90C06
1 Introduction
The main purpose of this paper is to develop an abstract framework for randomized subspace correction methods in convex optimization. While subspace correction methods [54, 58] generalize a broad class of iterative algorithms, convex optimization itself encompasses a wide range of applications, including nonlinear partial differential equations (PDEs), imaging, and data science. Consequently, the proposed framework is flexible and applicable to a very broad range of problems. Moreover, it accommodates highly general settings, such as weaker smoothness and convexity assumptions [40, 49].
Subspace correction methods [58] follow a divide-and-conquer strategy by decomposing the original problem into local subproblems defined on subspaces, which are solved independently. Many classical and modern iterative methods, including block relaxation, domain decomposition, and multigrid methods, can be viewed as instances of subspace correction methods. The theory has evolved over the past decades, covering both linear [58, 59] and nonlinear problems [11, 38, 54].
Block coordinate descent methods are prominent examples of subspace correction methods for convex optimization. These methods solve local subproblems restricted to blocks of coordinates, often via gradient or proximal steps. More general updates, like upper bound minimization [22, 46], are also possible. Their computational efficiency has led to widespread adoption. Key early results include [56, 61], and a comprehensive survey appears in [57]. Recent advances cover convergence of cyclic [6, 50], randomized [35, 47], accelerated [32, 33], and parallel variants [17, 34, 48]. Applications include deep neural network training [62, 63].
In numerical analysis, domain decomposition and multigrid methods are essential examples of subspace correction methods. Their convergence for smooth convex optimization was studied in [53, 54], and later extended to constrained and nonsmooth cases [1, 2, 3, 38]. These methods have been applied to various nonlinear variational problems, including PDEs [14, 30, 42], variational inequalities [2, 3, 41], elastoplasticity [11], and mathematical imaging [13, 21, 29].
Subspace correction methods are classified as parallel or successive depending on the order of subproblem updates [55, 58]. In parallel methods (additive Schwarz), all subproblems are solved concurrently; in successive methods (multiplicative Schwarz), subproblems are solved sequentially.
Randomizing the order of subproblem updates leads to randomized subspace correction methods [19, 23], which often exhibit better performance compared to fixed-order approaches. In particular, improved worst-case convergence rates under randomization are established in [23]. For quadratic optimization, a notable result in [51] shows that the worst-case complexities of cyclic coordinate descent methods and randomized coordinate descent methods are $\mathcal{O}(J^{4}\kappa\log(1/\epsilon))$ and $\mathcal{O}(J^{2}\kappa\log(1/\epsilon))$, respectively, where $J$ denotes the dimension of the variable, $\epsilon$ is a prescribed tolerance for the energy error, and $\kappa$ is a problem-dependent factor. More precisely, by constructing a specific quadratic objective, it was shown that both complexity bounds can be attained simultaneously. This demonstrates that cyclic coordinate descent methods can be up to $\mathcal{O}(J^{2})$ times worse than randomized coordinate descent methods, thereby clarifying the role of randomization: when the optimal ordering of subspaces is unknown, randomization can yield better performance than certain fixed-order strategies. One may also refer to [12] for an analysis of how randomization averages antisymmetric terms in descent inequalities. This mechanism is beneficial for the development of accelerated methods [32, 33] and has motivated extensive research on randomized methods in convex optimization [34, 36, 47].
This paper introduces an abstract framework for randomized subspace correction methods for convex composite optimization [36] on reflexive Banach spaces, accommodating diverse levels of smoothness and convexity [40, 49]. The framework unifies a wide range of decomposition strategies, including block partitioning [6, 35, 47] and overlapping domain decompositions commonly used in the numerical solution of PDEs [30, 42, 55, 58]. It supports both exact and inexact local solvers, encompassing methods such as coordinate descent, Bregman descent [16, 18, 24], and constraint decomposition [10, 52]. We establish convergence theorems that extend recent results, including those in [19, 47]. More precisely, by identifying a structural relationship between randomized and parallel subspace correction methods (see [23] for the linear case), we analyze the convergence of randomized subspace correction methods by leveraging existing analyses of parallel methods, such as [31, 38, 44]. Furthermore, we provide new analyses that yield sharper estimates under stronger, yet commonly adopted, assumptions; see, for example, [47].
The remainder of this paper is organized as follows. In Section 2, we present an abstract framework of randomized subspace correction methods for convex optimization. In Section 3, we derive convergence theorems under various conditions on the target problem. In Section 4, we provide explanations on how the proposed framework is related to existing results. In Section 5, we summarize possible applications of the proposed framework from diverse fields of science and engineering. Finally in Section 6, we conclude the paper with some remarks.
2 Subspace correction methods
This section presents an abstract framework for randomized subspace correction methods for convex optimization. In particular, we show that the convergence analysis of randomized subspace correction methods for convex optimization can be carried out within the framework of parallel subspace correction methods [38], extending the analogy previously established for linear problems [19, 23]. The proposed framework is highly versatile, accommodating diverse space decomposition settings for the model problem, a broad range of smoothness and convexity levels in the objective functional, and various types of inexact local solvers.
Let $V$ be a reflexive Banach space equipped with the norm $\|\cdot\|$. Its topological dual is denoted by $V^{*}$, and the duality pairing between $V^{*}$ and $V$ is written as
$\langle p,v\rangle=p(v),\quad p\in V^{*},\ v\in V.$
Throughout this paper, we adopt the convention $0/0=0$ for arguments of $\sup$ and $0/0=\infty$ for arguments of $\inf$.
2.1 Space decomposition and subspace correction
We consider the following abstract convex optimization problem:
(1) $\min_{v\in V}\left\{E(v):=F(v)+G(v)\right\},$
where $F\colon V\to\mathbb{R}$ is a Gâteaux differentiable and convex functional, and $G\colon V\to\overline{\mathbb{R}}$ is a proper, convex, and lower semicontinuous functional. The problem (1) is referred to as a composite optimization problem [36], as it involves a nonsmooth term $G$ in addition to the smooth term $F$. We further assume that the energy functional $E$ is coercive, which guarantees the existence of a minimizer $u\in V$ for the problem (1).
We assume that the solution space $V$ of (1) admits a space decomposition of the form
(2) $V=\sum_{j=1}^{J}V_{j},$
where each $V_{j}$, $j\in[J]=\{1,2,\dots,J\}$, is a closed subspace of $V$. The space decomposition (2) covers various algorithms, including block coordinate descent methods [35, 47], domain decomposition methods [38, 55], and multigrid methods [54, 60]. It is well known [59, Equation (2.15)] that the space decomposition (2) satisfies the stable decomposition property. Namely, for any $q\in[1,\infty)$, we have
(3) $\sup_{\|w\|=1}\inf_{\sum_{j=1}^{J}w_{j}=w}\left(\sum_{j=1}^{J}\|w_{j}\|^{q}\right)^{\frac{1}{q}}<\infty,$
where $w\in V$ and $w_{j}\in V_{j}$.
Subspace correction methods involve solving local problems defined on subspaces $\{V_{j}\}_{j=1}^{J}$. For a given $v\in V$, the optimal residual in a subspace $V_{j}$ is obtained by solving the local minimization problem
(4) $\min_{w_{j}\in V_{j}}E(v+w_{j}).$
Alternating minimization methods [5, 6] and certain domain decomposition methods (see, e.g., [2, 30, 42, 54]) fall into the category of subspace correction methods with exact local solvers as in (4). In contrast, block coordinate descent methods typically solve the local problem (4) inexactly, often using a single iteration of gradient descent [6, 35], proximal descent [33, 47], or Bregman descent [16, 18, 24]. Some methods further employ surrogate techniques, where (4) is replaced by an approximate problem with lower computational complexity; see, e.g., [10, 14, 52].
To encompass all these methods, following [38, 44], we consider local problems of the form
(5) $\min_{w_{j}\in V_{j}}\left\{E_{j}(w_{j};v):=F_{j}(w_{j};v)+G_{j}(w_{j};v)\right\},$
where $F_{j}(\cdot;v)\colon V_{j}\to\mathbb{R}$ and $G_{j}(\cdot;v)\colon V_{j}\to\overline{\mathbb{R}}$ are convex functionals for each $v\in V$. The functionals $F_{j}(\cdot;v)$ and $G_{j}(\cdot;v)$ serve as approximations to the exact local functionals $F(v+\cdot)$ and $G(v+\cdot)$ on $V$, respectively. An example of (5) corresponding to a proximal descent step is presented in Example 2.1. Additional examples can be found in [38, Section 6.4].
Example 2.1.
If we set
$F_{j}(w_{j};v)=F(v)+\langle F^{\prime}(v),w_{j}\rangle+\frac{1}{2\tau_{j}}\|w_{j}\|^{2},\ G_{j}(w_{j};v)=G(v+w_{j}),\quad v\in V,\ w_{j}\in V_{j},$
for some $\tau_{j}>0$, then the local problem (5) corresponds to a single proximal descent step [33, 47] with step size $\tau_{j}$ for minimizing $E(v+w_{j})$.
The abstract parallel subspace correction method for solving the convex optimization problem (1), based on the space decomposition (2) and local solvers (5), is presented in Algorithm 1.
Algorithm 1 Parallel subspace correction method for (1)
 Given the step size $\tau>0$:
 Choose $u^{(0)}\in\operatorname{dom}G$.
for $n=0,1,2,\dots$ do
  for $j\in[J]$ in parallel do
   $w_{j}^{(n+1)}\in\operatornamewithlimits{\arg\min}_{w_{j}\in V_{j}}E_{j}(w_{j};u^{(n)})$
  end for
  $u^{(n+1)}=u^{(n)}+\tau\sum_{j=1}^{J}w_{j}^{(n+1)}$
end for
Another type of subspace correction method is the successive subspace correction method, in which the local problems in the subspaces are solved sequentially. In this paper, we focus on a particular variant known as the randomized subspace correction method, where the order of the local problems is chosen randomly; see Algorithm 2.
Algorithm 2 Randomized subspace correction method for (1)
 Choose $u^{(0)}\in\operatorname{dom}G$.
for $n=0,1,2,\dots$ do
  Sample $j\in[J]$ from the uniform distribution on $[J]$.
  $w_{j}^{(n+1)}\in\operatornamewithlimits{\arg\min}_{w_{j}\in V_{j}}E_{j}(w_{j};u^{(n)})$
  $u^{(n+1)}=u^{(n)}+w_{j}^{(n+1)}$
end for
Remark 2.2.
The randomized subspace correction method presented in Algorithm 2 can also be generalized to the case of nonuniform sampling, as considered in, e.g., [35, 45, 48]. For brevity, we do not discuss this in detail in this paper.
2.2 Descent property
In what follows, we denote by $d$ and $d_{j}$ the Bregman divergences associated with $F$ and $F_{j}$, respectively:
$d(w;v)=F(v+w)-F(v)-\langle F^{\prime}(v),w\rangle,\quad$ $v,w\in V,$
$d_{j}(w_{j};v)=F_{j}(w_{j};v)-F_{j}(0;v)-\langle F_{j}^{\prime}(0;v),w_{j}\rangle,\quad$ $v\in V,\ w_{j}\in V_{j}.$
To ensure the convergence of the randomized subspace correction method, we adopt the assumptions on the local problem (5) summarized in Section 2.2. We note that Section 2.2 provides a more general framework than several recent works, as it extends the smooth settings in [31, 44] to the nonsmooth case, and employs a broader local stability assumption (see Section 2.2(c)) than the one used in [38].
{assumption}
[local problems] For any $j\in[J]$ and $v\in V$, the local functionals $F_{j}(\cdot;v)\colon V_{j}\to\mathbb{R}$ and $G_{j}(\cdot;v)\colon V_{j}\to\overline{\mathbb{R}}$ satisfy the following:
  1. (a)
    (convexity) The functional $F_{j}(\cdot;v)$ is Gâteaux differentiable and convex, while $G_{j}(\cdot;v)$ is proper, convex, and lower semicontinuous. Moreover, the composite functional $E_{j}(\cdot;v)$ is coercive.
  2. (b)
    (consistency) We have
    $F_{j}(0;v)=F(v),\quad G_{j}(0;v)=G(v),$
    and
    $\langle F_{j}^{\prime}(0;v),w_{j}\rangle=\langle F^{\prime}(v),w_{j}\rangle,\quad w_{j}\in V_{j}.$
  3. (c)
    (stability) For some $\omega\in(0,1]\cup(1,\rho)$, we have
    $d(w_{j};v)\leq\omega d_{j}(w_{j};v),\quad G(v+w_{j})\leq G_{j}(w_{j};v),\quad w_{j}\in V_{j},$
    where the constant $\rho$ is defined as
    (6) $\rho=\min_{j\in[J]}\inf_{v\in V,\ w_{j}\in V_{j}}\frac{\langle d_{j}^{\prime}(w_{j};v),w_{j}\rangle}{d_{j}(w_{j};v)}.$
The constant $\rho$ defined in (6) is always greater than or equal to $1$ as a consequence of Section 2.2(a,b). In the case of linear problems, one can verify that $\rho=2$ [44, Example 1], which is consistent with [55, 59]. A nonlinear example where $\rho>1$ is provided in [31, Example A.2]. In Lemma 2.3, which is a nonsmooth extension of [44, Lemma 1], we show that Section 2.2 ensures that solving each local problem leads to a decrease in the global energy.
Lemma 2.3.
For $j\in[J]$ and $v\in V$, let
(7) $\hat{w}_{j}\in\operatornamewithlimits{\arg\min}_{w_{j}\in V_{j}}E_{j}(w_{j};v).$
Under Section 2.2, we have
$E(v)-E(v+\hat{w}_{j})\geq\left(1-\frac{\omega}{\rho}\right)\langle d_{j}^{\prime}(\hat{w}_{j};v),\hat{w}_{j}\rangle\geq 0.$
Proof 2.4.
The optimality condition for $\hat{w}_{j}$ reads as
(8) $G_{j}(w_{j};v)-G_{j}(\hat{w}_{j};v)\geq\langle F_{j}^{\prime}(\hat{w}_{j};v),\hat{w}_{j}-w_{j}\rangle,\quad w_{j}\in V_{j}.$
In particular, for $w_{j}=0$, we obtain
[…]
where the first inequality follows from Section 2.2(b,c). On the other hand, by Section 2.2(c), we have
[…]
Summing (9) and (10) completes the proof.
As a corollary, we deduce that the energy in Algorithm 2 decreases monotonically; see Corollary 2.5.
Corollary 2.5.
Suppose that Section 2.2 holds. In the randomized subspace correction method (Algorithm 2), the sequence $\{E(u^{(n)})\}$ is decreasing.
In Lemma 2.6, we present a refined version (cf. [44, Lemma 2]) of the generalized additive Schwarz lemma for the composite optimization problem (1), originally introduced in [38, Lemma 4.5].
Lemma 2.6.
Suppose that Section 2.2(a,b) holds. For $v\in V$, we have
(11) $\hat{w}:=\sum_{j=1}^{J}\hat{w}_{j}\in\operatornamewithlimits{\arg\min}_{w\in V}\left\{\langle F^{\prime}(v),w\rangle+\inf_{w=\sum_{j=1}^{J}w_{j}}\sum_{j=1}^{J}(d_{j}+G_{j})(w_{j};v)\right\},$
where $\hat{w}_{j}$, $j\in[J]$ was given in (7). Moreover, we have
(12) $\inf_{w=\sum_{j=1}^{J}}\sum_{j=1}^{J}(d_{j}+G_{j})(w_{j};v)=\sum_{j=1}^{J}(d_{j}+G_{j})(\hat{w}_{j};v).$
Proof 2.7.
We closely follow the argument in the proof of [31, Lemma 4.2]. Throughout the proof, we define
$d(w;v)=\inf_{\sum_{j=1}^{J}w_{j}=w}(d_{j}+G_{j})(w_{j};v),\quad w\in V.$
Let $w\in V$ be arbitrary. For any $w_{j}\in V_{j}$, $j\in[J]$, such that $w=\sum_{j=1}^{J}w_{j}$, we have
[…]
Here, the first inequality and the last equality follow from the fact that $\hat{w}=\sum_{j=1}^{J}\hat{w}_{j}\in\sum_{j=1}^{J}V_{j}$. By minimizing the last line of (13) over all decompositions $(w_{j})_{j=1}^{J}$, we obtain
(14) $\langle F^{\prime}(v),\hat{w}\rangle+d(\hat{w};v)\leq\sum_{j=1}^{J}\left(\langle F^{\prime}(v),\hat{w}_{j}\rangle+(d_{j}+G_{j})(\hat{w}_{j};v)\right)\leq\langle F^{\prime}(v),w\rangle+d(w;v),$
which implies (11). Finally, setting $w=\hat{w}$ in (14) yields (12).
Lemma 2.6 shows that, to analyze the convergence rate of the parallel subspace correction method (Algorithm 1), it suffices to estimate the following quantity [31, 44]:
(15) $\Psi(u^{(n)}):=\min_{w\in V}\left\{\langle F^{\prime}(u^{(n)}),w\rangle+\inf_{w=\sum_{j=1}^{J}w_{j}}\sum_{j=1}^{J}(d_{j}+G_{j})(w_{j};u^{(n)})\right\}-JG(u^{(n)}).$
In the randomized subspace correction method (Algorithm 2), the update at each iteration is determined by a randomly chosen subspace. Consequently, the total energy after one step, $E(u^{(n+1)})$, is a random variable depending on the sampling of the index $j\in[J]$. To analyze its expected descent behavior, we consider the conditional expectation of the energy $\mathbb{E}[E(u^{(n+1)})|u^{(n)}]$, which represents the expected value of the energy at the next iteration given the current iterate $u^{(n)}$. In Theorem 2.8, we show that the conditional expectation $\mathbb{E}[E(u^{(n+1)})|u^{(n)}]$ can be estimated using (15), indicating that its analysis can proceed along similar lines as that of the parallel method.
Theorem 2.8.
Suppose that Section 2.2 holds. In the randomized subspace correction method (Algorithm 2), we have
$\mathbb{E}[E(u^{(n+1)})|u^{(n)}]\leq E(u^{(n)})+\frac{\theta}{J}\Psi(u^{(n)}),\quad n\geq 0,$
where $\Psi(u^{(n)})$ was given in (15), and the constant $\theta$ is given by
(16) $\theta=\begin{cases}1,\quad&\text{ if }\omega\in(0,1],\\ \frac{\rho-\omega}{\rho-1},\quad&\text{ if }\omega\in(1,\rho).\end{cases}$
Proof 2.9.
Fix any $n\geq 0$. For each $j\in[J]$, let $w_{j}^{(n+1)}\in V_{j}$ be a minimizer of $F_{j}(w_{j};u^{(n)})$. We first consider the case $\omega\in(0,1]$. It follows that
(17) $\begin{split}\mathbb{E}[E(u^{(n+1)})\mid u^{(n)}]&=\frac{1}{J}\sum_{j=1}^{J}E(u^{(n)}+w_{j}^{(n+1)})\\ &\stackrel{{\scriptstyle\text{(i)}}}{{\leq}}F(u^{(n)})+\frac{1}{J}\sum_{j=1}^{J}\left[\langle F^{\prime}(u^{(n)}),w_{j}^{(n+1)}\rangle+(d_{j}+G_{j})(w_{j}^{(n+1)};u^{(n)})\right]\\ &\stackrel{{\scriptstyle\text{(ii)}}}{{=}}E(u^{(n)})+\frac{1}{J}\Psi(u^{(n)}),\end{split}$
which is the desired result. Here, (i) follows from Section 2.2(c), and (ii) follows from Lemma 2.6.
Now consider the case $\omega\in(1,\rho)$. Proceeding similarly as in (17), we obtain
(18) $\mathbb{E}[E(u^{(n+1)})\mid u^{(n)}]\leq E(u^{(n)})+\frac{1}{J}\Psi(u^{(n)})+\frac{\omega-1}{J}\sum_{j=1}^{J}d_{j}(w_{j}^{(n+1)};u^{(n)}).$
Meanwhile, from (6) and Lemma 2.3, it follows that
(19) $\begin{split}\frac{1}{J}\sum_{j=1}^{J}d_{j}(w_{j}^{(n+1)};u^{(n)})&\leq\frac{1}{J\rho}\sum_{j=1}^{J}\langle d_{j}^{\prime}(w_{j}^{(n+1)};u^{(n)}),w_{j}^{(n+1)}\rangle\\ &\leq\frac{1}{\rho-\omega}\left(E(u^{(n)})-\mathbb{E}[E(u^{(n+1)})\mid u^{(n)}]\right).\end{split}$
Combining (18) and (19) yields the desired result.
3 Convergence theorems
In this section, we present convergence theorems for the randomized subspace correction method (Algorithm 2) under various conditions on the energy functional $E$. The convergence results are derived by invoking Theorem 2.8 and following arguments similar to those developed for the parallel subspace correction method (Algorithm 1), as presented in [31, 38].
Given the initial iterate $u^{(0)}\in\operatorname{dom}G$ of Algorithm 2, we define
(20) $K_{0}=\{v\in V:E(v)\leq E(u^{(0)})\},\quad R_{0}=\sup_{v\in K_{0}}\|v-u\|.$
The convexity and coercivity of $E$ imply that $K_{0}$ is bounded and convex, and in particular, $R_{0}<\infty$. Moreover, by Corollary 2.5, the sequence $\{u^{(n)}\}$ generated by Algorithm 2 remains entirely within $K_{0}$.
Theorem 2.8 implies that, to estimate the convergence rate of Algorithm 2, it suffices to estimate $\Psi(u^{(n)})$ defined in (15). From the expression of $\Psi(u^{(n)})$, the following stable decomposition assumption arises naturally (cf. [38, Assumption 4.1]).
{assumption}
[stable decomposition] For some $q>1$, the following holds: for any bounded convex subset $K$ of $V$, we have
(21a) $C_{K}:=q\sup_{v,v+w\in K}\inf\frac{\sum_{j=1}^{J}d_{j}(w_{j};v)}{\|w\|^{q}}<\infty,$
where the infimum is taken over $w_{j}\in V_{j}$, $j\in[J]$, satisfying
(21b) $w=\sum_{j=1}^{J}w_{j},\quad\sum_{j=1}^{J}G_{j}(w_{j};v)\leq G(v+w)+(J-1)G(v).$
Examples of stable decompositions satisfying Section 3 will be provided in Sections 4 and 5; see also [38, Section 6]. A notable observation made in [44, Lemma 3] is that, in the case of smooth problems, i.e., when $G=0$ in (1) and $G_{j}=0$, $j\in[J]$, in (5), Section 3 need not be assumed, but instead holds automatically under a mild smoothness condition on each $d_{j}$; see Proposition 3.1.
Proposition 3.1.
In the case of smooth problems, i.e., when $G=0$ in (1) and $G_{j}=0$, $j\in[J]$, in (5), suppose that Section 2.2 holds. Furthermore, assume that for some $q>1$, the following holds: for any bounded convex subsets $K\subset V$ and $K_{j}\subset V_{j}$ with $0\in K_{j}$, we have
$\sup_{v\in K,\ w_{j}\in K_{j}}\frac{d_{j}(w_{j};v)}{\|w_{j}\|^{q}}<\infty.$
Then we have
$C_{K}=q\sup_{v,\,v+w\in K}\inf_{w=\sum_{j=1}^{J}w_{j}}\frac{\sum_{j=1}^{J}d_{j}(w_{j};v)}{\|w\|^{q}}<\infty.$
Proof 3.2.
This result is a special case of [31, Lemma 4.10], which relies on the stable decomposition property (3).
The following lemma provides a preliminary estimate for $\Psi(u^{(n)})$ under Section 3. Although the proof follows a similar argument to that in [38, Appendix A.3], we include it here for completeness.
Lemma 3.3.
Suppose that Section 3 holds. Then we have
(22) $\Psi(u^{(n)})\leq\min_{t\in[0,1]}\bigg\{t\langle F^{\prime}(u^{(n)}),u-u^{(n)}\rangle+\frac{C_{K_{0}}}{q}t^{q}\|u-u^{(n)}\|^{q}\\ +G((1-t)u^{(n)}+tu)\bigg\}-G(u^{(n)}),\quad n\geq 0,$
where $\Psi(u^{(n)})$, $K_{0}$, and $C_{K_{0}}$ were given in (15), (21), and (20), respectively.
Proof 3.4.
From the definition (15) of $\Psi(u^{(n)})$, we have
(23) $\begin{split}\Psi(u^{(n)})&=\min_{w\in V}\left\{\langle F^{\prime}(u^{(n)}),w\rangle+\inf_{w=\sum_{j=1}^{J}w_{j}}\sum_{j=1}^{J}(d_{j}+G_{j})(w_{j};u^{(n)})\right\}-JG(u^{(n)})\\ &\leq\min_{u^{(n)}+w\in K_{0}}\left\{\langle F^{\prime}(u^{(n)}),w\rangle+\inf_{w=\sum_{j=1}^{J}w_{j}}\sum_{j=1}^{J}(d_{j}+G_{j})(w_{j};u^{(n)})\right\}-JG(u^{(n)})\\ &\leq\min_{u^{(n)}+w\in K_{0}}\left\{\langle F^{\prime}(u^{(n)}),w\rangle+\frac{C_{K_{0}}}{q}\|w\|^{q}+G(u^{(n)}+w)\right\}-G(u^{(n)}),\end{split}$
where the last inequality follows from Section 3. The proof is complete upon replacing $w$ in the last line of (23) with $t(u-u^{(n)})$ for some $t\in[0,1]$.
If one can derive suitable bounds for the term
(24) $\tilde{\Psi}(t;u^{(n)}):=t\langle F^{\prime}(u^{(n)}),u-u^{(n)}\rangle+G\big((1-t)u^{(n)}+tu\big)-G(u^{(n)}),\quad t\in[0,1],$
then, in view of Lemma 3.3, one can proceed to obtain quantitative convergence bounds. Sharper estimates for $\tilde{\Psi}(t;u^{(n)})$ can be established under stronger assumptions on $F$ and $G$. Accordingly, we consider several such cases in the following subsections.
Remark 3.5.
An improved estimate for $\Psi(u^{(n)})$ compared to that in Lemma 3.3 can be obtained under a stronger assumption than Section 3. Suppose that the following global stable decomposition condition holds:
(25) $C_{V}:=q\sup_{v,\,v+w\in V}\inf\frac{\sum_{j=1}^{J}d_{j}(w_{j};v)}{\|w\|^{q}}<\infty,$
where the infimum is taken over (21b). Under this global condition, the constraint $u^{(n)}+w\in K_{0}$ in (23) is no longer needed, and consequently, the restriction $t\in[0,1]$ in (22) can be relaxed to $t\geq 0$. This improvement will be useful later in our analysis; see Remarks 3.12 and 3.15.
3.1 General problems
Without imposing additional assumptions on $F$ and $G$, we can still obtain the following upper bound for (24) using the convexity of $F$ and $G$:
(26) $\tilde{\Psi}(t;u^{(n)})\leq-t(E(u^{(n)})-E(u)).$
By combining Theorem 2.8, Lemma 3.3, and (26), we obtain the following convergence theorem for the randomized subspace correction method (Algorithm 2).
Theorem 3.6.
Suppose that Sections 2.2 and 3 hold. In the randomized subspace correction method (Algorithm 2), if $\zeta_{0}:=E(u^{(0)})-E(u)>C_{K_{0}}R_{0}^{q}$, then
$\mathbb{E}[E(u^{(1)})]-E(u)\leq\left(1-\frac{\theta}{J}\left(1-\frac{1}{q}\right)\right)\zeta_{0},$
where $\theta$, $K_{0}$, $R_{0}$, and $C_{K_{0}}$ were given in (16), (20), and (21). Otherwise, we have
$\mathbb{E}[E(u^{(n)})]-E(u)\leq\frac{C}{(n+(C/\zeta_{0})^{1/\beta})^{\beta}},\quad n\geq 0,$
where
$\beta=q-1,\quad C=\left(\frac{Jq}{\theta}\right)^{q-1}C_{K_{0}}R_{0}^{q}.$
Proof 3.7.
We write $\zeta_{n}=E(u^{(n)})-E(u)$. Combining Lemma 3.3 and (26) yields
(27) $\Psi(u^{(n)})\leq\min_{t\in[0,1]}\left\{-t\zeta_{n}+\frac{t^{q}C_{K_{0}}}{q}\|u-u^{(n)}\|^{q}\right\}.$
Applying the argument in [31, Equation (B.2)], we obtain
(28) $\Psi(u^{(n)})\leq\begin{cases}-\left(1-\frac{1}{q}\right)\zeta_{n}&\text{ if }\zeta_{n}>C_{K_{0}}R_{0}^{q},\\ -\left(1-\frac{1}{q}\right)\frac{\zeta_{n}^{\frac{q}{q-1}}}{(C_{K_{0}}R_{0}^{q})^{\frac{1}{q-1}}}&\text{ if }\zeta_{n}\leq C_{K_{0}}R_{0}^{q}.\end{cases}$
Combining Theorem 2.8 and (28), we obtain
(29) $\mathbb{E}[E(u^{(n+1)})|u^{(n)}]-E(u)\leq\begin{cases}\left(1-\frac{\theta}{J}\left(1-\frac{1}{q}\right)\right)\zeta_{n}&\text{ if }\zeta_{n}>C_{K_{0}}R_{0}^{q},\\ \zeta_{n}-\frac{\theta}{J}\left(1-\frac{1}{q}\right)\frac{\zeta_{n}^{\frac{q}{q-1}}}{(C_{K_{0}}R_{0}^{q})^{\frac{1}{q-1}}}&\text{ if }\zeta_{n}\leq C_{K_{0}}R_{0}^{q}.\end{cases}$
This proves the desired result for the case $\zeta_{0}>C_{K_{0}}R_{0}^{q}$. On the other hand, by Corollary 2.5, the condition $\zeta_{0}\leq C_{K_{0}}R_{0}^{q}$ ensures $\zeta_{n}\leq C_{K_{0}}R_{0}^{q}$. By the law of total expectation and Jensen inequality
$\mathbb{E}[\zeta_{n}]^{\frac{q}{q-1}}\leq\mathbb{E}[\zeta_{n}^{\frac{q}{q-1}}],$
we obtain
$\mathbb{E}[\zeta_{n+1}]\leq\mathbb{E}[\zeta_{n}]-\frac{\theta}{J}\left(1-\frac{1}{q}\right)\frac{\mathbb{E}[\zeta_{n}]^{\frac{q}{q-1}}}{(C_{K_{0}}R_{0}^{q})^{\frac{1}{q-1}}}$
if $\zeta_{0}\leq C_{K_{0}}R_{0}^{q}$. Finally, invoking [31, Lemma B.2] completes the proof for the case $\zeta_{0}\leq C_{K_{0}}R_{0}^{q}$.
The generality of the assumptions in Theorem 3.6 enables a broad range of applications, particularly in scenarios where $F$ exhibits a weaker level of smoothness than the standard smoothness condition [40]; see [31, 37] for concrete examples.
Remark 3.8.
Since Algorithm 2 is expected to visit all subspaces $\{V_{j}\}_{j=1}^{J}$ on average within $J$ iterations, it is natural to examine the convergence behavior of Algorithm 2 at iteration counts that are integer multiples of $J$, say $nJ$. We observe that the expected energy error has an upper bound independent of $J$:
$\mathbb{E}\bigl[E(u^{(nJ)})\bigr]-E(u)\leq\frac{\hat{C}}{(n+(\hat{C}/\zeta_{0})^{1/\beta})^{\beta}},\quad n\geq 0,$
where
$\beta=q-1,\qquad\hat{C}=\left(\frac{q}{\theta}\right)^{q-1}C_{K_{0}}R_{0}^{q}.$
3.2 Sharp problems
Meanwhile, in many applications, the energy functional $F$ satisfies the sharpness condition [49], summarized in Section 3.2, which is also known as the Hölderian error bound or the Łojasiewicz inequality [8, 61].
{assumption}
[sharpness] For some $p>1$, the function $F$ satisfies the following: for any bounded convex subset $K$ of $V$ satisfying $u\in K$, we have
(30) $\mu_{K}:=p\inf_{v\in K}\frac{F(v)-F(u)}{\|v-u\|^{p}}>0.$
If we additionally assume that Section 3.2 holds, then we can derive the following improved convergence theorem for the randomized subspace correction method.
Theorem 3.9.
Suppose that Sections 2.2, 3, and 3.2 hold. In the randomized subspace correction method (Algorithm 2), we have the following:
  1. (a)
    In the case $p=q$, we have
    $\mathbb{E}[E(u^{(n)})]-E(u)\leq\left(1-\frac{\theta}{J}\left(1-\frac{1}{q}\right)\min\left\{1,\frac{\mu_{K_{0}}}{qC_{K_{0}}}\right\}^{\frac{1}{q-1}}\right)^{n}\zeta_{0},\quad n\geq 0,$
    where $\zeta_{0}=E(u^{(0)})-E(u)$, and $\theta$, $K_{0}$, $C_{K_{0}}$, and $\mu_{K_{0}}$ were given in (16), (20), (21) and (30), respectively.
  2. (b)
    In the case $p>q$, if $\zeta_{0}>\left(\frac{p}{\mu_{K_{0}}}\right)^{\frac{q}{p-q}}C_{K_{0}}^{\frac{p}{p-q}}$, then we have
    $\mathbb{E}[E(u^{(1)})]-E(u)\leq\left(1-\frac{\theta}{J}\left(1-\frac{1}{q}\right)\right)\zeta_{0}.$
    Otherwise, we have
    $\mathbb{E}[E(u^{(n)})]-E(u)\leq\frac{C}{\left(n+(C/\zeta_{0})^{1/\beta}\right)^{\beta}},\quad n\geq 0,$
    where
    $\beta=\frac{p(q-1)}{p-q},\quad C=\left(\frac{Jpq}{(p-q)\theta}\right)^{\frac{p(q-1)}{p-q}}\left(\frac{p}{\mu_{K_{0}}}\right)^{\frac{q}{p-q}}C_{K_{0}}^{\frac{p}{p-q}}.$
Proof 3.10.
We again set $\zeta_{n}=E(u^{(n)})-E(u)$. By Section 3.2 and (27), we obtain
(31) $\Psi(u^{(n)})\leq\min_{t\in[0,1]}\left\{-t\zeta_{n}+\frac{t^{q}C_{K_{0}}}{q}\|u-u^{(n)}\|^{q}\right\}\leq\min_{t\in[0,1]}\left\{-t\zeta_{n}+\frac{t^{q}p^{\frac{q}{p}}C_{K_{0}}}{q\mu_{K_{0}}^{\frac{q}{p}}}\zeta_{n}^{\frac{q}{p}}\right\}.$
We first consider the case $p=q$. It follows from (31) that
(32) $\Psi(u^{(n)})\leq\min_{t\in[0,1]}\left\{-t\zeta_{n}+\frac{t^{q}qC_{K_{0}}}{q\mu_{K_{0}}}\zeta_{n}\right\}\leq-\zeta_{n}\left(1-\frac{1}{q}\right)\min\left\{1,\frac{\mu_{K_{0}}}{qC_{K_{0}}}\right\}^{\frac{1}{q-1}}.$
Combining Theorem 2.8 and (32), we obtain
$\mathbb{E}[E(u^{(n+1)})|u^{(n)}]-E(u)\leq\left(1-\frac{\theta}{J}\left(1-\frac{1}{q}\right)\min\left\{1,\frac{\mu_{K_{0}}}{qC_{K_{0}}}\right\}^{\frac{1}{q-1}}\right)\zeta_{n}.$
Invoking the law of total expectation yields the desired result.
Next, we consider the case $p>q$. Applying the argument in [31, Equation (B.5)] to (31), we obtain
(33) $\Psi(u^{(n)})\leq\begin{cases}-\left(1-\frac{1}{q}\right)\zeta_{n}&\text{ if }\zeta_{n}>\left(\frac{p}{\mu_{K_{0}}}\right)^{\frac{q}{p-q}}C_{K_{0}}^{\frac{p}{p-q}},\\[2.15277pt] -\left(1-\frac{1}{q}\right)\left(\frac{\mu_{K_{0}}}{p}\right)^{\frac{q}{p(q-1)}}\frac{\zeta_{n}^{\frac{q(p-1)}{p(q-1)}}}{C_{K_{0}}^{\frac{1}{q-1}}}&\text{ if }\zeta_{n}\leq\left(\frac{p}{\mu_{K_{0}}}\right)^{\frac{q}{p-q}}C_{K_{0}}^{\frac{p}{p-q}}.\end{cases}$
Combining Theorem 2.8 and (33), we derive
(34) $\mathbb{E}[E(u^{(n+1)})|u^{(n)}]\leq\begin{cases}\left(1-\frac{\theta}{J}\left(1-\frac{1}{q}\right)\right)\zeta_{n}&\text{ if }\zeta_{n}>\left(\frac{p}{\mu_{K_{0}}}\right)^{\frac{q}{p-q}}C_{K_{0}}^{\frac{p}{p-q}},\\[2.15277pt] \zeta_{n}-\frac{\theta}{J}\left(1-\frac{1}{q}\right)\left(\frac{\mu_{K_{0}}}{p}\right)^{\frac{q}{p(q-1)}}\frac{\zeta_{n}^{\frac{q(p-1)}{p(q-1)}}}{C_{K_{0}}^{\frac{1}{q-1}}}&\text{ if }\zeta_{n}\leq\left(\frac{p}{\mu_{K_{0}}}\right)^{\frac{q}{p-q}}C_{K_{0}}^{\frac{p}{p-q}}.\end{cases}$
Observing that (29) and (34) share the same structure, proceeding as in the proof of Theorem 3.6 completes the proof.
Remark 3.11.
Similar to Remark 3.8, the expected energy error at the $nJ$th iteration of Algorithm 2 admits an upper bound that is independent of $J$. In the case $p=q$, we have
$\mathbb{E}[E(u^{(nJ)})]-E(u)\leq\left(1-\frac{\theta}{J}\left(1-\frac{1}{q}\right)\min\left\{1,\frac{\mu_{K_{0}}}{qC_{K_{0}}}\right\}^{\frac{1}{q-1}}\right)^{nJ}\zeta_{0}\\ \leq\exp\left(-n\theta\left(1-\frac{1}{q}\right)\min\left\{1,\frac{\mu_{K_{0}}}{qC_{K_{0}}}\right\}^{\frac{1}{q-1}}\right)\zeta_{0}.$
The case $p>q$ can be analyzed analogously to Remark 3.8. Namely, we have
$\mathbb{E}\bigl[E(u^{(nJ)})\bigr]-E(u)\leq\frac{\hat{C}}{(n+(\hat{C}/\zeta_{0})^{1/\beta})^{\beta}},\quad n\geq 0,$
where
$\beta=\frac{p(q-1)}{p-q},\quad\hat{C}=\left(\frac{pq}{(p-q)\theta}\right)^{\frac{p(q-1)}{p-q}}\left(\frac{p}{\mu_{K_{0}}}\right)^{\frac{q}{p-q}}C_{K_{0}}^{\frac{p}{p-q}}.$
Remark 3.12.
In the case of smooth problems, i.e., when $G=0$ in (1) and $G_{j}=0$, $j\in[J]$ in (5), the global stable decomposition condition (25), together with the argument in Remark 3.5, yields the following simplified estimate corresponding to Theorem 3.9(a):
$\mathbb{E}\big[F(u^{(n)})\big]-F(u)\leq\left(1-\frac{\theta}{J}\left(1-\frac{1}{q}\right)\left(\frac{\mu_{K_{0}}}{qC_{V}}\right)^{\frac{1}{q-1}}\right)^{n}\zeta_{0},\quad n\geq 0.$
3.3 Strongly convex problems
Next, we consider a stronger condition than Section 3.2, stated in Section 3.3. Recall that a functional $H\colon V\to\overline{\mathbb{R}}$ is said to be $\mu$-strongly convex on a convex set $K\subset V$ if
$H((1-t)v+tw)\leq(1-t)H(v)+tH(w)-t(1-t)\frac{\mu}{2}\|v-w\|^{2},\quad t\in[0,1],\ v,w\in K.$
Note that Section 3.3 implies Section 3.2 with $p=2$.
{assumption}
[strong convexity] For any bounded convex subset $K$ of $V$ satisfying $u\in K$, $E$ and $F$ are $\mu_{K}$- and $\mu_{F,K}$-strongly convex on $K$, respectively, for some $\mu_{K}>0$ and $\mu_{F,K}\geq 0$.
Note that $\mu_{K}\geq\mu_{F,K}$. Under Section 3.3, we have the following upper bound for (24):
(35) $\begin{split}\tilde{\Psi}(t;u^{(n)})&\leq E((1-t)u^{(n)}+tu)-E(u^{(n)})-\frac{\mu_{F,K_{0}}}{2}t^{2}\|u-u^{(n)}\|^{2}\\ &\leq-t(E(u^{(n)})-E(u))+\left(\frac{\mu_{K_{0}}-\mu_{F,K_{0}}}{2}t^{2}-\frac{\mu_{K_{0}}}{2}t\right)\|u-u^{(n)}\|^{2},\end{split}$
where the first and second inequalities follow from the strong convexity of $F$ and $E$, respectively.
As Section 3.3 imposes a stronger condition than Section 3.2, the convergence rate established in Theorem 3.9(b) is guaranteed under Section 3.3. However, by using (35), we can derive an even sharper estimate for the convergence rate, as presented in Theorem 3.13, when $q=2$ in Section 3. A similar result appears in [42, Theorem 3.3].
Theorem 3.13.
Suppose that Sections 2.2, 3, and 3.3 hold with $q=2$. In the randomized subspace correction method (Algorithm 2), we have
$\mathbb{E}[E(u^{(n)})]-E(u)\leq\left(1-\frac{\theta}{J}\min\left\{1,\frac{\mu_{K_{0}}}{C_{K_{0}}+\mu_{K_{0}}-\mu_{F,K_{0}}}\right\}\right)^{n}\zeta_{0},\quad n\geq 0,$
where $\zeta_{0}=F(u^{(0)})-F(u)$, and $\theta$, $K_{0}$, $C_{K_{0}}$, $\mu_{K_{0}}$, and $\mu_{F,K_{0}}$ were given in (16), (20), (21) and Section 3.3.
Proof 3.14.
Note that $C_{K_{0}}+\mu_{K_{0}}-\mu_{F,K_{0}}$ is positive under Section 3.3. Combining Lemma 3.3 and (35) yields
$\Psi(u^{(n)})\leq\min_{t\in[0,1]}\left\{-t(E(u^{(n)})-E(u))+\left(\frac{C_{K_{0}}+\mu_{K_{0}}-\mu_{F,K_{0}}}{2}t^{2}-\frac{\mu_{K_{0}}}{2}t\right)\|u-u^{(n)}\|^{2}\right\}.$
If $\frac{\mu_{K_{0}}}{C_{K_{0}}+\mu_{K_{0}}-\mu_{F,K_{0}}}\leq 1$, then setting $t=\frac{\mu_{K_{0}}}{C_{K_{0}}+\mu_{K_{0}}-\mu_{F,K_{0}}}$ in (36) gives
(37) $\Psi(u^{(n)})\leq-\frac{\mu_{K_{0}}}{C_{K_{0}}+\mu_{K_{0}}-\mu_{F,K_{0}}}(E(u^{(n)})-E(u)).$
Otherwise, if $\frac{\mu_{K_{0}}}{C_{K_{0}}+\mu_{K_{0}}-\mu_{F,K_{0}}}>1$, then setting $t=1$ in (36) yields
(38) $\begin{split}\Psi(u^{(n)})&\leq-(E(u^{(n)})-E(u))+\frac{C_{K_{0}}-\mu_{F,K_{0}}}{2}\|u-u^{(n)}\|^{2}\\ &\leq\left(\frac{C_{K_{0}}-\mu_{F,K_{0}}}{\mu_{K_{0}}}-1\right)(E(u^{(n)})-E(u))\\ &\leq-(E(u^{(n)})-E(u)),\end{split}$
where the second inequality follows again from Section 3.3. Combining Theorem 2.8 with (37) and (38) completes the proof.
Remark 3.15.
Similar to Remark 3.12, in the case of smooth problems, i.e., when $G=0$ in (1) and $G_{j}=0$, $j\in[J]$ in (5), the linear convergence rate in Theorem 3.13 simplifies under the global stable decomposition condition (25) as follows:
$\mathbb{E}[F(u^{(n)})]-F(u)\leq\left(1-\frac{\theta}{J}\frac{\mu_{K_{0}}}{C_{V}+\mu_{K_{0}}-\mu_{F,K_{0}}}\right)^{n}\zeta_{0},\quad n\geq 0.$
4 Derivation of related methods
To highlight the versatility of the randomized subspace correction framework presented in this paper, we demonstrate in this section how several related results can be derived from it, including the randomized subspace correction method for linear problems [19, 23], block coordinate descent methods [33, 35, 47], and operator splitting methods [25].
Throughout this section, we assume that $V$ is a Hilbert space equipped with an inner product $(\cdot,\cdot)$ and the induced norm $\|\cdot\|$. Moreover, we identify $V$ with its topological dual space $V^{*}$ (cf. [60]).
4.1 Linear problems
The randomized subspace correction method for linear problems has been previously studied in [19, 23]. Here, we demonstrate how the general framework introduced in this paper recovers these existing results.
In (1), we set
$F(v)=\frac{1}{2}(Av,v)-(f,v),\quad G(v)=0,\quad v\in V,$
where $A\colon V\to V$ is a linear operator induced by a continuous, symmetric, and coercive bilinear form on $V$, and $f\in V$. Then, it is readily observed that (1) reduces to the linear problem
(39) $Au=f.$
In the local problem (5), for each $j\in[J]$, we set
(40) $F_{j}(w_{j};v)=F(v)+(F^{\prime}(v),w_{j})+\frac{1}{2}(R_{j}^{-1}w_{j},w_{j}),\quad G_{j}(w_{j};v)=0,\quad w_{j}\in V_{j},\ v\in V,$
where $R_{j}\colon V_{j}\to V_{j}$ is a linear operator induced by a continuous, symmetric, and coercive bilinear form on $V_{j}$. Then, the solution $\hat{w}_{j}$ of (5) is given by
$\hat{w}_{j}=-R_{j}Q_{j}(Av-f),$
where $Q_{j}\colon V\to V_{j}$ denotes the orthogonal projection onto $V_{j}$. Hence, we observe that Algorithm 2 reduces to the randomized subspace correction method for linear problems introduced in [19, 23].
To analyze the algorithm, it suffices to verify Sections 2.2, 3, and 3.3. We adopt the $\|\cdot\|_{A}$-norm defined by $\|\cdot\|_{A}:=(A\cdot,\cdot)^{\frac{1}{2}}$. Note that Section 2.2(a, b) are trivially satisfied due to (40). In Section 2.2(c), we have $\rho=2$ [44, Example 1], and the assumption reduces to the following condition: for some $\omega\in(0,2)$, we require
(41) $(Aw_{j},w_{j})\leq\omega(R_{j}^{-1}w_{j},w_{j}),\quad w_{j}\in V_{j},$
which corresponds to the standard assumption on local solvers, as found in, e.g., [55, Assumption 2.4] and [59, Equation (4.6)]. Moreover, Proposition 3.1 ensures that Section 3 holds. Specifically, we have $q=2$, and for any bounded convex subset $K$ of $V$, it follows that
$C_{K}\leq\sup_{w\in V}\inf_{w=\sum_{j=1}^{J}w_{j}}\frac{\sum_{j=1}^{J}(R_{j}^{-1}w_{j},w_{j})}{(Aw,w)}=\lambda_{\min}(T)^{-1},$
where the operator $T\colon V\to V$ is given by
(42) $T=\sum_{j=1}^{J}R_{j}Q_{j}A,$
and the last equality follows from the well-known estimate in [59, Equation (2.17)]; see also [38, Section 4.1]. Finally, we observe that Section 3.3 holds with
$\mu_{K}=\mu_{F,K}\geq 1$
for any bounded convex subset $K$ of $V$. Therefore, under the condition (41), the convergence estimate in Remark 3.15 yields
$\mathbb{E}[\|u^{(n)}-u\|_{A}^{2}]\leq\left(1-\frac{\theta\lambda_{\min}(T)}{J}\right)^{n}\|u^{(0)}-u\|_{A}^{2},\quad n\geq 0,$
where $\theta$ and $T$ were given in (16) and (42), respectively (cf. [19, Theorem 1(b)]).
4.2 Block coordinate descent methods
Block coordinate descent methods are important instances of subspace correction methods for convex optimization. We discuss how randomized block coordinate descent methods [33, 35, 47] can be interpreted within the framework of the randomized subspace correction method; see also [25, Section 3.3] and [38, Section 6.4] for related discussions.
We begin by presenting the standard setting for block coordinate descent methods [33, 47]. Assume that the space $V$ is given by the direct sum of subspaces $V_{j}$, $j\in[J]$:
$V=\bigoplus_{j=1}^{J}V_{j},$
so that each $v\in V$ can be represented in block form as [48, Proposition 1]
(43) $v=(v_{1},v_{2},\dots,v_{J}),\quad v_{j}\in V_{j}.$
In the composite optimization problem (1), we assume that $F$ is block smooth. That is, there exist positive constants $L_{j}$, $j\in[J]$, such that
(44) $F(v+w_{j})\leq F(v)+(F^{\prime}(v),w_{j})+\frac{L_{j}}{2}\|w_{j}\|^{2},\quad v\in V,\ w_{j}\in V_{j}.$
In addition, we assume that the functional $G$ is block separable, meaning it admits the decomposition
(45) $G(v)=\sum_{j=1}^{J}G^{j}(v_{j}),\quad v\in V,$
for some functionals $G^{j}\colon V_{j}\to\overline{\mathbb{R}}$.
In the local problem (5), for each $j\in[J]$, we set
(46) $\begin{aligned} F_{j}(w_{j};v)&=F(v)+(F^{\prime}(v),w_{j})+\frac{L_{j}}{2}\|w_{j}\|^{2},\\ G_{j}(w_{j};v)&=G(v+w_{j}),\end{aligned}\quad w_{j}\in V_{j},\ v\in V.$
Then, it is readily seen that the local problem (5) computes the increment of a single proximal descent step with step size $1/L_{j}$ applied to the $j$th block coordinate; see Example 2.1. Consequently, Algorithm 2 reduces to the randomized block coordinate descent method considered in [33, 47].
Next, we demonstrate that our convergence theory recovers the results established in [47] for the randomized block coordinate descent method. It is straightforward to verify that Section 2.2 holds with $\omega=1$, based on (44), (45), and (46). Moreover, using (45), we obtain (cf. [38, Section 6.4])
$\sum_{j=1}^{J}G_{j}(w_{j};v)=G(v+w)+(J-1)G(v),\quad v,w\in V.$
From (44) and (46), we have
$\sum_{j=1}^{J}d_{j}(w_{j};v)\leq\frac{1}{2}\|w\|_{L}^{2},\quad v,w\in V,$
where the norm $\|\cdot\|_{L}$ is defined by
$\|v\|_{L}=\left(\sum_{j=1}^{J}L_{j}\|v_{j}\|^{2}\right)^{\frac{1}{2}},\quad v\in V.$
That is, Section 3 holds with the $\|\cdot\|_{L}$-norm, $q=2$, and $C_{K}=1$ for any bounded convex subset $K\subset V$. Consequently, by Theorem 3.6, the randomized block coordinate descent method satisfies
$\mathbb{E}[E(u^{(n)})]-E(u)\leq\frac{2JR_{0}^{2}}{n+2JR_{0}^{2}/(E(u^{(0)})-E(u))},\quad n\geq 0,$
provided that $E(u^{(0)})-E(u)$ is sufficiently small. Otherwise, the method exhibits linear convergence. Here, $R_{0}$ is defined by (20) with respect to the $\|\cdot\|_{L}$-norm. This result is consistent with [47, Theorem 5] and [33, Equation (16)].
Now suppose further that $F$ and $G$ are $\mu_{F}$- and $\mu_{G}$-strongly convex with respect to the $\|\cdot\|_{L}$-norm, respectively, for some $\mu_{F},\mu_{G}\geq 0$ with $\mu_{F}+\mu_{G}>0$. In this case, Theorem 3.13 implies the following linear convergence rate:
$\mathbb{E}[E(u^{(n)})]-E(u)\leq\left(1-\frac{1}{J}\frac{\mu_{F}+\mu_{G}}{1+\mu_{G}}\right)^{n}(E(u^{(0)})-E(u)),\quad n\geq 0,$
which agrees with [47, Theorem 7].
4.3 Operator splitting methods
By utilizing the duality between subspace correction and operator splitting methods developed in [25], we can derive a randomized operator splitting method from the randomized subspace correction method.
As a model problem for operator splitting methods, we consider the following optimization problem involving the sum of multiple convex functionals:
(47) $\min_{v\in V}\left\{F(v)+\sum_{j=1}^{J}G_{j}(B_{j}v)\right\},$
where $V$ and each $W_{j}$, $j\in[J]$, are Hilbert spaces, $B_{j}\colon V\to W_{j}$ are continuous linear operators, and $F\colon V\to\mathbb{R}$ and $G_{j}\colon W_{j}\to\overline{\mathbb{R}}$ are proper, convex, and lower semicontinuous functionals. For simplicity, we assume that $F$ is strongly convex and smooth so that $F^{\prime}$ is invertible [4].
A randomized Peaceman–Rachford-type splitting method for solving (47) is presented in Algorithm 3. At each iteration of Algorithm 3, an index $j\in[J]$ is selected at random, and an optimization problem involving only the functional $G_{j}$ is solved. Recall that $d$ denotes the Bregman divergence associated with $F$.
Algorithm 3 Randomized Peaceman–Rachford splitting algorithm for (47)
 Choose $u^{(0)}\in V$ and $v_{1}^{(0)},\dots,v_{J}^{(0)}\in V$.
for $n=0,1,2,\dots$ do
  Sample $j\in[J]$ from the uniform distribution on $[J]$.
  $u^{(n+1)}=\operatornamewithlimits{\arg\min}_{v\in V}\left\{d(v-u^{(n)};u^{(n)})+(v_{j}^{(n)},v)+G_{j}(B_{j}v)\right\}$
  $v_{i}^{(n+1)}=\begin{cases}v_{i}^{(n)}+F^{\prime}(u^{(n+1)})-F^{\prime}(u^{(n)}),&\quad\text{ if }i=j,\\ v_{i}^{(n)},&\quad\text{ if }i\neq j,\end{cases}\quad i\in[J]$
end for
As summarized in Theorem 4.1, and similar to [25, Theorem 5.4], we can deduce that Algorithm 3 is a dualization of the randomized subspace correction method applied to the following dual problem:
(48) $\min_{(p_{j})_{j=1}^{J}\in\bigoplus_{j=1}^{J}W_{j}}\left\{F^{*}\left(-\sum_{j=1}^{J}B_{j}^{*}p_{j}\right)+\sum_{j=1}^{J}G_{j}^{*}(p_{j})\right\},$
where $F^{*}\colon V\to\mathbb{R}$ and $G_{j}^{*}\colon W_{j}\to\overline{\mathbb{R}}$ denote the Legendre–Fenchel conjugates of $F$ and $G_{j}$, respectively, and $B_{j}^{*}$ denotes the adjoint of $B_{j}$. Note that (48) is an instance of (1) in which the nonsmooth part is block separable (cf. (45)).
Theorem 4.1.
Let $\{u^{(n)}\}$, $\{(v_{j}^{(n)})_{j=1}^{J}\}$, and $\{(p_{j}^{(n)})_{j=1}^{J}\}$ be the sequences generated by the randomized Peaceman–Rachford splitting algorithm for solving (47) (Algorithm 3) and the randomized subspace correction method with exact local problems for solving (48). If
$u^{(0)}=(F^{*})^{\prime}\left(-\sum_{j=1}^{J}B_{j}^{*}p_{j}^{(0)}\right),\quad v_{j}^{(0)}=-B_{j}^{*}p_{j}^{(0)},\quad j\in[J],$
then we have
$u^{(n)}\stackrel{{\scriptstyle\text{a.s.}}}{{=}}(F^{*})^{\prime}\left(-\sum_{j=1}^{J}B_{j}^{*}p_{j}^{(n)}\right),\quad v_{j}^{(n)}\stackrel{{\scriptstyle\text{a.s.}}}{{=}}-B_{j}^{*}p_{j}^{(n)},\quad j\in[J],\ n\geq 1.$
5 Applications
In this section, we present applications of the randomized subspace correction method to a range of problems arising in diverse areas of science and engineering.
5.1 Linear problems
A fundamental class of problems to which the randomized subspace correction method applies is linear problems [19, 23], as discussed in Section 4.1. Given the extensive literature on subspace correction methods for linear systems, particularly those arising from the numerical discretization of elliptic PDEs, we omit detailed discussion for brevity. We refer the reader to [20, 43, 55] and the references therein.
5.2 Nonlinear partial differential equations
The randomized subspace correction method is also applicable to nonlinear PDEs that admit convex variational formulations. As an example, we briefly present the $s$-Laplacian problem, which was also considered in [30, 54]. For other examples of nonlinear PDEs, one may refer to, e.g., [14, 42].
We consider the following nonlinear problem:
$\begin{split}-\operatorname{div}(|\nabla u|^{s-2}\nabla u)=f&\quad\text{ in }\Omega,\\ u=0&\quad\text{ on }\partial\Omega,\end{split}$
where $\Omega\subset\mathbb{R}^{2}$ is a bounded polygonal domain, $s\in(1,\infty)$, and $f\in W^{-1,s^{*}}(\Omega)$ with $1/s+1/s^{*}=1$. It is well-known that the above problem admits a weak formulation given by the following convex optimization problem:
(49) $\min_{v\in W_{0}^{1,s}(\Omega)}\left\{\frac{1}{s}\int_{\Omega}|\nabla v|^{s}\,dx-\langle f,v\rangle\right\}.$
To numerically solve (49), we employ a finite element discretization. Let $\mathcal{T}_{h}$ be a quasi-uniform triangulation of $\Omega$, where $h$ denotes the characteristic element diameter. We denote by $S_{h}(\Omega)$ the lowest-order Lagrangian finite element space defined on $\mathcal{T}_{h}$, incorporating the homogeneous essential boundary condition. Then, the finite element approximation of (49) is given by
(50) $\min_{v\in S_{h}(\Omega)}\left\{\frac{1}{s}\int_{\Omega}|\nabla v|^{s}\,dx-\langle f,v\rangle\right\}.$
We observe that (50) is an instance of the abstract problem (1). Namely, (50) corresponds to (1) with
$V=S_{h}(\Omega),\quad F(v)=\frac{1}{s}\int_{\Omega}|\nabla v|^{s}\,dx-\langle f,v\rangle,\quad G(v)=0.$
For the space decomposition (2), two-level overlapping domain decomposition and multigrid decomposition with exact local solvers were studied in [30, 38] and [54], respectively. In both cases, one can verify that Sections 3 and 3.2 hold with $p=\max\{s,2\}$ and $q=\min\{s,2\}$ with respect to the $W^{1,s}$-norm, allowing the application of Theorem 3.9 to establish convergence of the method.
5.3 Variational inequalities
Another important class of problems is variational inequalities, which find applications in computational mechanics and optimal control. As an illustrative example, we consider a second-order problem [3, 52]; see [2, 11, 41] for further examples.
We consider the following variational inequality: find $u\in K$ such that
$\int_{\Omega}\nabla u\cdot\nabla v\,dx-\langle f,v\rangle\geq 0,\quad v\in K,$
where $f\in H^{-1}(\Omega)$, and $K$ is a subset of $H_{0}^{1}(\Omega)$ representing a pointwise inequality constraint:
$K=\{v\in H_{0}^{1}(\Omega):v\leq g\text{ a.e. in }\Omega\},$
for some $g\in C(\Omega)$. This problem admits the equivalent optimization formulation
$\min_{v\in K}\left\{\frac{1}{2}\int_{\Omega}|\nabla v|^{2}\,dx-\langle f,v\rangle\right\}.$
Using a finite element discretization defined on $S_{h}(\Omega)$, we obtain the discrete problem
(51) $\min_{v\in S_{h}(\Omega)}\left\{\frac{1}{2}\int_{\Omega}|\nabla v|^{2}\,dx-\langle f,v\rangle+\chi_{K_{h}}(v)\right\},$
where $K_{h}=K\cap S_{h}(\Omega)$, and $\chi_{K_{h}}$ denotes the indicator functional of $K_{h}$, defined as $\chi_{K_{h}}(v)=0$ if $v\in K_{h}$ and $\chi_{K_{h}}(v)=\infty$ otherwise. We observe that (51) is an instance of the general form (1). Specifically, we obtain (51) if we set
$V=S_{h}(\Omega),\quad F(v)=\frac{1}{2}\int_{\Omega}|\nabla v|^{2}\,dx-\langle f,v\rangle,\quad G(v)=\chi_{K_{h}}(v).$
Two-level overlapping domain decomposition and multigrid decomposition with exact local solvers were studied in [1, 3]. In addition, constraint decomposition methods, based on localized constraints and interpretable as instances of inexact local solvers within our framework [38, Section 6.4], were considered in [52]. In all cases, one can verify that Sections 3 and 3.3 hold with $q=2$ with respect to the $H^{1}$-norm, thereby enabling the application of Theorem 3.13 to analyze the convergence of the algorithms.
5.4 Total variation minimization
Total variation minimization is a fundamental problem in mathematical imaging; see, e.g., [12]. Given a bounded polygonal domain $\Omega\subset\mathbb{R}^{2}$, we consider the variational problem
(52) $\min_{v\in BV(\Omega)}\left\{\frac{1}{2}\int_{\Omega}(v-f)^{2}\,dx+TV(v)\right\},$
where $TV(v)$ denotes the total variation of $v$, $BV(\Omega)$ is the space of functions of bounded variation, and $f\in L^{2}(\Omega)$.
Designing subspace correction methods for solving (52) is particularly challenging due to the nonseparable structure of the total variation term [39]. Indeed, it was shown in [27] that standard domain decomposition methods generally fail to satisfy the stable decomposition condition stated in Section 3.
One viable approach is to instead consider the dual formulation of (52[26], which reads:
(53) $\min_{\mathbf{p}\in H_{0}(\operatorname{div};\Omega)}\frac{1}{2}\int_{\Omega}(\operatorname{div}\mathbf{p}+f)^{2}\,dx\quad\text{subject to}\quad\|\mathbf{p}\|_{L^{\infty}(\Omega)}\leq 1.$
By employing the Raviart–Thomas finite element discretization introduced in [28], we obtain the discrete problem
(54) $\min_{\mathbf{p}\in\mathbf{S}_{h}(\Omega)}\left\{\frac{1}{2}\int_{\Omega}(\operatorname{div}\mathbf{p}+f)^{2}\,dx+\chi_{\mathbf{K}_{h}}(\mathbf{p})\right\},$
where $\mathbf{S}_{h}(\Omega)$ denotes the lowest-order Raviart–Thomas finite element space on the mesh $\mathcal{T}_{h}$, and $\mathbf{K}_{h}\subset\mathbf{S}_{h}(\Omega)$ is a convex set encoding the constraint $\|\mathbf{p}\|_{L^{\infty}(\Omega)}\leq 1$. The discrete problem (54) fits the abstract formulation (1), with the following identifications:
$V=\mathbf{S}_{h}(\Omega),\quad F(\mathbf{p})=\frac{1}{2}\int_{\Omega}(\operatorname{div}\mathbf{p}+f)^{2}\,dx,\quad G(\mathbf{p})=\chi_{\mathbf{K}_{h}}(\mathbf{p}).$
Schwarz-type domain decomposition methods for the dual problem (54) were analyzed in [39], and related constraint decomposition techniques were introduced in [13]. In both cases, it can be verified that Section 3 holds with $q=2$, allowing Theorem 3.6 to be invoked for convergence analysis. Similar results can also be established for finite difference discretizations of (53); see [21, 29].
Meanwhile, to tackle the primal problem (52) directly, one may employ the operator splitting approach described in Section 4.3 to design domain decomposition methods. In this context, it was shown in [25, Section 5.5.1] that the resulting methods coincide with those proposed in [27].
5.5 Multinomial logistic regression
Multinomial logistic regression is a fundamental tool for classification problems in machine learning. Given a labeled dataset $\{(x_{j},y_{j})\}_{j=1}^{J}\subset\mathbb{R}^{d}\times[k]$, where each $x_{j}\in\mathbb{R}^{d}$ represents a data point and $y_{j}\in[k]$ its corresponding class label, the regression model is formulated as follows:
(55) $\min_{\theta\in\mathbb{R}^{(d+1)k}}\left\{\frac{1}{J}\sum_{j=1}^{J}\left(\operatorname{LSE}_{k}(X_{j}^{T}\theta)-\hat{x}^{T}\theta\right)+\frac{\alpha}{2}\|\theta\|^{2}\right\},$
where $\theta=[w_{1}^{T},b_{1},\dots,w_{k}^{T},b_{k}]^{T}\in\mathbb{R}^{(d+1)k}$ is the parameter vector, $\operatorname{LSE}_{k}$ denotes the log-sum-exp function over $k$ classes, $X_{j}\in\mathbb{R}^{(d+1)k\times k}$ is defined as $X_{j}=I_{k}\otimes[x_{j}^{T},1]^{T}$, and $\alpha$ is a positive regularization hyperparameter. The vector $\hat{x}\in\mathbb{R}^{(d+1)k}$ is a constant vector depending only on the dataset; see [25, Example 2.10] for details. Note that Eq. 55 is an instance of (47).
In big data settings, where the number of data points $J$ is very large, operator splitting methods such as stochastic gradient descent [7], which process individual data points or mini-batches, are widely used to solve (55). The randomized Peaceman–Rachford splitting algorithm presented in Algorithm 3 is one such method.
Thanks to Theorem 4.1, the convergence analysis of Algorithm 3 for solving (55) reduces to analyzing the randomized subspace correction method for solving the following dual problem [25]:
(56) $\min_{(p_{j})_{j=1}^{J}\in\mathbb{R}^{Jk}}\left\{\frac{1}{2J\alpha}\left\|\sum_{j=1}^{J}X_{j}p_{j}-\hat{x}\right\|^{2}+\sum_{j=1}^{J}\operatorname{LSE}_{k}^{*}(p_{j})\right\}.$
Note that the term $\sum_{j=1}^{J}\operatorname{LSE}_{k}^{*}(p_{j})$ is block separable. Therefore, by an argument analogous to that in Section 4.2, one can verify that Section 3 holds with $q=2$. Consequently, we may apply Theorem 3.6 to establish the convergence of the randomized subspace correction method for solving (56).
6 Conclusion
In this paper, we introduced an abstract framework for randomized subspace correction methods for convex optimization. We established convergence theorems that are applicable to a broad range of scenarios involving space decomposition, local solvers, and varying levels of problem regularity. Furthermore, we demonstrated that these theorems unify and extend several relevant recent results.
This work opens several avenues for future research. One important direction is the development of convergence theory for cyclic successive subspace correction methods. This is particularly relevant for multigrid methods [1, 54], where the hierarchical structure of subspaces plays a critical role. We note that a sharp convergence analysis of cyclic methods for linear problems can be found in [9, 59]. The block coordinate descent method with $q=2$ was analyzed in [6]; however, the results there are not sharp in the sense that the analysis does not recover the sharp estimate established for the linear case.
Another promising direction is to extend the proposed framework to nonconvex problems. Recent work [15] has shown that randomized block coordinate descent methods are effective for a certain class of nonconvex problems. Generalizing the randomized subspace correction framework to accommodate such settings remains an interesting and challenging problem.